OpenStudy (anonymous):

y''+9y=1, y(0)=0, y'(pi/2)=3 must be solve by green's function

OpenStudy (abb0t):

You could use undetermined coefficients to solve this. Do you know about undetermined coefficients?

OpenStudy (anonymous):

i found c1cos3x+c2sin3x but i didn't know how to find c1 and c2

OpenStudy (abb0t):

To find \(C_1\) and \(C_2\), simply plug in the initial conditions. Then find the derivative of the solution. And again apply the initial conditions. You sould get a system.

OpenStudy (abb0t):

Remember: \(y=y_c+y_p\), you already found you're complimentary solution: \(y_c\), no wyou need to find \(y_p\)

OpenStudy (anonymous):

yc=c1cos3x+c2sin3x yc'=-3c1sin3x+3c2cos3x y(0)=0 c1=0 y'(pi/2)=3 3c1=3 c1=1????? couldn't find c2

OpenStudy (anonymous):

this is a BVP problem and i need to find yc only

OpenStudy (anonymous):

y(0)=0 c1=0 then let c2=1 so y1=sin3x is this right?

OpenStudy (anonymous):

y'(pi/2)=3 3c1=3 c1=1.......i don't know how to find y2

OpenStudy (abb0t):

Wait, hold on, let me do this really quick im getting a paper.

OpenStudy (anonymous):

actually the question need to solve by green's function so i gotta find y1 and y2 first

OpenStudy (abb0t):

i still think you need to get your particular solution. The solution i got was: \(y = c_1\cos(3x)+c_2\sin(3x)+\frac{ 1 }{ 9 }\)

OpenStudy (abb0t):

Plugging in the intial conditions: \[c_1+\frac{ 1 }{ 9 } = 0\]

OpenStudy (abb0t):

Actually, I'm getting lost in this myself now. I will respond to you. loL. Give me a minute to figure this one out. Lol

OpenStudy (anonymous):

i have to use this formula to solve this question

OpenStudy (anonymous):

How is \(G\) relevant?

OpenStudy (anonymous):

\[yp(x)=\int\limits_{xo}^{x} G(x,t)f(t)dt\]

OpenStudy (anonymous):

\[yp(x)=\int\limits_{a}^{x}\frac{ y1(t)y2(x) }{ w(t)}f(t)dt+\int\limits_{x}^{b}\frac{ y1(x)y2(t) }{w(t)}f(t)dt\]

OpenStudy (anonymous):

I would love to help, but I'm extremely confused.

OpenStudy (anonymous):

this BVP question gotta be solve by green's function